a) (i) Use EDTA structure to describe hexadentate chelating agent (2mks)
(ii) State four types of EDTA titrations (2mks)

Consider the titration of 50mL Mg 2+ sample in a pH = 10.00 buffer with 0.05M EDTA as titrant. Calculate the amount of Mg in ppm if the titre is 5ml of EDTA (3mks)
Calculate the pH of a 1000-mL buffer solution containing 0.200M NH4OH and 0.300M NH4CI. (3mks)
1
Expert's answer
2020-04-08T13:37:56-0400

a) i EDTA or ethylenediaminetetraacetic acid is a novel molecule for complexing metal ions. It is a polyprotic acid containing four carboxylic acid groups (acidic hydrogens are red) and two amine groups with lone pair electrons (green dots). The classic structural formula is given below.

ii Direct titrationthe cations are titrated directly with standard solution EDTA using eriochrome black T as the indicator. Back titration: a known excess of standard solution EDTA is added to the solution containing the analyte.

0.005 x 0.05 = 0.00025

120 ppm.

II. pH = pKa + lg(aNH4OH / aNH4Cl) = 9.25 + lg(0.2 / 0.3) = 8.67